# Simulation-based Approach for Fast Optimal Control of a Stefan Problem with Application to Cell Therapy

### v1

@@ 4.2.1 Problem description and formulation
Besides using the fastest thawing as in Problem 1, some past studies considered cases in which the freezing and thawing rates are controlled, i.e., the rate of temperature change is kept constant (Seki and Mazur, 2008; Jang et al., 2017; Baboo et al., 2019; Bojic et al., 2021). As such, this problem focuses on controlling the rate of change in the average temperature during thawing by manipulating the heater temperature. The definition of the average temperature is as given in (10).
For a fixed heater temperature, the rate of temperature change is not constant (Figs.  2DE). Thus, the heater temperature needs to be manipulated. Consider the optimal control problem
@@
where $d\Theta_{\textrm{avg}}/d\tau$ is the rate of change in the average temperature and $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ is the target value (setpoint) set to 0.04 for demonstration in this case study. For the optimization-based approaches, the control vector is parameterized using a piecewise linear function, with the number of control intervals $n_{c}=16$ (Appendix A).
@@ 4.2.2 DAE Reformulation
For the simulation-based technique, first analyze the objective function in (17). This objective function is minimized when the rate of change in average temperature is equal to the setpoint. Replacing the objective function in (17) with the algebraic equation $d\Theta_{\textrm{avg}}/d\tau=(d\Theta_{\textrm{avg}}/d\tau)_{\textrm{sp}}$ results in the system of DAEs
@@
where $\Theta_{b}$ is now treated as a state. The differential index of (18) is 2, which can be solved using GEKKO as explained in Section 3.3.
@@ 4.2.3 Solution comparison
where $(d\Theta_{\textrm{avg}}/d\tau)_{k}$ is the rate of change in average temperature resulting from solving (17) or (18) evaluated at each time point $k$ in the time span [$\tau_{0},\tau_{f}$] and $n_{k}$ is the number of sampling time points. Small $||e||_{2}$ corresponds to the rate of change in average temperature being close to the target value, indicating that the heater temperature is optimal; i.e., the solution method is accurate.
The optimal solutions to Problem 2 obtained from the optimization- and simulation-based approaches are illustrated in Fig.  3. Most solution techniques predict the same heater temperature profile except the methods with fmincon that predict a slightly higher temperature (Fig.  3A). As time progresses, the heater temperature increases to compensate for a reduction in the temperature difference (driving force) between the heater and product. The rate of change in average temperature can be controlled nearly perfectly at 0.04 for IPOPT, CasADi, GEKKO, and the simulation-based method, while this value fluctuates up to about 0.046 for fmincon, implying that fmincon fails to converge to the optimal solution in this case (Fig.  3B). It is also evident that the average temperature increases linearly at the rate of 0.04 in all cases except for fmincon (Fig.  3C).
Figure 3: Trajectories for the optimal (A) heater temperature, (B) rate of change in average temperature, and (C) average temperature simulated via the optimization- and simulation-based approaches in Problem 2.
@@
The simulation-based approach is by far the fastest solution method, accelerating the computation by 40$\times$ to 90$\times$ compared to the typical solvers IPOPT and fmincon (Table 2). Without parallel computing, optimization with IPOPT is fastest, followed by GEKKO, fmincon, and CasADi. Parallel computing can significantly reduce the computation time for fmincon. With the direct multiple shooting method (opt_mCasADi), the number of states resulting from spatial discretization is large. For example, with $n=20$, there are 41 states for each control interval and 656 states for $n_{c}=16$. This large nonlinear optimization is computationally expensive to solve compared to other approaches. In terms of accuracy, fmincon produces the largest error, which is about 30–40 fold higher than for the other techniques. This indicates that the solution obtained from fmincon is not optimal, agreeing with the plots shown in Fig.  3. Although both fmincon and IPOPT employ the interior point methods, the detailed implementations (e.g., gradient/hessian calculation, scaling) are not exactly the same, which could result in different accuracy and performance (see detailed discussion and comparison of these solvers in Rojas-Labanda and Stolpe (2015)). The simulation-based and other optimization-based methods except fmincon have similar accuracy. Despite an increase in the differential index, the simulation-based approach is still much faster than the optimization-based techniques while giving the same level of accuracy.
@@ 4.3.1 Problem description and formulation
One of the primary interests in a Stefan problem is the moving solid-liquid interface. The literature has studied and shown the importance of interface position tracking and optimization in various applications (Hill and Wu, 1994; Brezina et al., 2018; Srisuma et al., 2023b, for example). This problem focuses on the control of the interface velocity, i.e., the evolution of a melting/freezing process.
@@
where $dS/d\tau$ is the interface velocity and $\left(dS/d\tau\right)_{\mathrm{sp}}$ is the target velocity set to $-0.1$ for demonstration in this case study. In Problem 1 (Fig.  2BC), the velocity of the moving interface is not constant if the heater temperature is fixed. In this case study, the aim is to control the velocity of the moving interface to be constant at $(dS/d\tau)_{\mathrm{sp}}$ by manipulating the heater temperature throughout the process as formulated in (20). For the optimization-based approaches, the control vector is parameterized using a piecewise linear function, with the number of control intervals $n_{c}=12$ (see Appendix A).
@@ 4.3.2 DAE Reformulation
To apply the simulation-based technique, a similar approach used in Problem 2 can be used, but here we instead enforce the moving interface velocity by modifying the interface equation (6) to be equal to $(dS/d\tau)_{\textrm{sp}}$. Consequently, (20) is reformulated as the DAE system
@@
The differential index of (21) is $n$, which is the number of grid points resulting from spatial discretization. Therefore, if the spatial discretization of the domain is made finer to increase numerical accuracy, the differential index of (21) increases. In this case, $n=20$, and hence (21) becomes an index-20 DAE system, making the problem even more complicated and difficult to solve compared to the index-2 system in Problem 2.
To explore the possibility of using the index reduction technique, MATLAB’s reduceDAEIndex was used to reduce the differential index of (21). However, the index reduction process was not complete even after 20 hours of simulation. Obviously, this technique is not feasible for our problem as mentioned before in Section 3.3. Therefore, GEKKO’s DAE solver is needed.
@@ 4.3.3 Solution comparison
The optimal solutions to Problem 3 are shown in Fig.  4. The optimal heater temperature predicted by each technique is nearly identical except that there is a large drop around the end of the process for the fmincon cases (Fig.  4A). The interface velocity can be controlled at about $-0.1$ all the time for IPOPT, CasADi, GEKKO, and the simulation-based technique, while some fluctuation is observed for fmincon (Fig.  4B). This indicates that fmincon is less accurate than the others approaches, which is a similar trend observed before in Problem 2. With the interface velocity controlled, the interface position recedes linearly at the rate of 0.1 (Fig.  4C).
Figure 4: Trajectories for the optimal (A) heater temperature, (B) interface velocity, and (C) interface position via the optimization- and simulation-based approaches in Problem 3.
For the computational performance, the simulation-based technique is about 60$\times$ to 900$\times$ faster than all of the optimization-based approaches (Table 3). IPOPT, GEKKO, and CasADi are comparable for the direct single shooting method, while the multiple shooting method is much slower. Parallel computing can significantly accelerate fmincon but is still much slower than the simulation-based approach. fmincon is the least accurate method; the error is higher than for the other approaches by more than an order of magnitude. Except for fmincon, the simulation-based approach has similar accuracy as the optimization-based approaches, with the error of about $2$$\times$$10^{-5}$.
@@
The simulation-based approach, however, has one limitation associated with the DAE reformulation. As described in Section 3.3, the DAE reformulation is guaranteed to be optimal when at least one of the constraints, bounds, or algebraic equations resulting from reformulating the objective function is active. This is typically the case when the objective function of the original optimal control problem can be minimized by solving each subproblem resulting from control vector parameterization individually – in other words, local minimization at every control/time interval (subproblem) is equivalent to global minimization of the whole process (original problem). For example, in Problem 1, if the thawing time of every control interval is minimized, the total thawing time is automatically minimized. In Problems 2 and 3, if the rate of temperature change and interface velocity are controlled locally for each control interval, the whole process is also controlled. These objective functions are common in many applications, for which the simulation-based technique can be applied, including lithium-ion batteries of various chemistries (Berliner et al., 2022; Galuppini et al., 2023; Matschek et al., 2023) and microwave lyophilization (Srisuma et al., 2023a). An example of a system in which the method is not optimal is a reactor with multiple chemical reactions, with the goal of maximizing the amount of desired product at the end of the process. Maximizing the amount of product during the first control interval could concurrently increase the amount of some byproducts that can degrade the desired product later. In this situation, local optimization is not equivalent to global optimization. As such, an algebraic equation/constraint may not be active, so an optimization-based technique is needed. Although the simulation-based approach does not give an optimal solution to such problems, it can be used to provide an initialization to the optimization solver.
@@ 4.4 Problem 4: Sensitivity analysis
In Problems 2 and 3, the simulation-based approach is tested with one value of $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ and $(dS/d\tau)_{\textrm{sp}}$, which are 0.04 and $-0.1$, respectively. To demonstrate the robustness of our framework, this section conducts a sensitivity analysis by varying $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ and $(dS/d\tau)_{\textrm{sp}}$ and employs the simulation-based approach to solve the problems.
The simulation-based approach provides accurate solutions irrespective of the values of $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ and $(dS/d\tau)_{\textrm{sp}}$ (Figs.  5 and 6). The rate of change in average temperature and interface velocity are at the setpoints except for $(dS/d\tau)_{\textrm{sp}}=-0.15$. An interface velocity of $-0.15$ is too fast for the given heater temperature, so the upper bound is active, and that the heater temperature cannot be increased further to achieve the target velocity. The target velocity that is larger (smaller in magnitude) than the peak of Fig.  2C, which about $-0.12$, will never violate the bound (see Appendix C for a formal proof).
For cases where the bound is active, a simulation switching technique can be used to transition between active constraints, resulting in a hybrid discrete/continuous dynamic simulation (Feehery and Barton, 1998; Barton et al., 2000; Berliner et al., 2022). First, initialize a DAE simulation as usual. Next, terminate the current simulation when the bound on any control vector is active and then initialize a new DAE simulation with that active bound/constraint. Perform this switching whenever there is a change in the active constraints, which results in a mixed continuous-discrete DAE system instead of a pure continuous system. The complexity of this procedure depends on the choice of a DAE solver. For example, MATLAB’s and Julia’s DAE solvers have a built-in function to track all variables and terminate a simulation when the specified condition (aka event) is met, which facilitates the implementation of such switching. The switching technique can also be used to handle cases where the control trajectory is discontinuous (see examples in Berliner et al. (2022); Srisuma et al. (2023a)). GEKKO’s DAE solvers, on the other hand, do not have a built-in option for handling switches, and so this process needs to be executed manually (see our software/code for implementation). This switching technique can also be used for path constraints.
Figure 5: Trajectories for the optimal (A) heater temperature, (B) rate of change in average temperature, and (C) average temperature simulated via the simulation-based approach at four different values of $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ in Problem 4.
Figure 6: Trajectories for the optimal (A) heater temperature, (B) interface velocity, and (C) interface position simulated via the simulation-based approach at four different values of $(dS/d\tau)_{\textrm{sp}}$ in Problem 4.
@@ 5 Conclusion
Our proposed framework was demonstrated and benchmarked against a variety of optimization-based approaches – IPOPT, fmincon, CasADi, and GEKKO – for case studies involving the optimal control of a Stefan problem for cell thawing. The case studies include the minimization of the thawing time, the control of the average temperature, and the control of the interface velocity, transformed into DAE systems of index 1, 2, and 20, respectively. At the same level of accuracy, the simulation-based approach is more than an order of magnitude faster than every optimization-based method in all cases. The simulation-based approach offers an efficient and reliable framework for solving the optimal control problems, serving as a promising alternative to the traditional techniques in applications where speed is crucial, e.g., real-time online model predictive control.
Some future directions of interest are to (1) generalize the approach and (2) improve the computational performance. The ultimate goal is a highly efficient optimal control solver applicable to problems with a variety of objective functions and constraints that arise in multiple fields. Implementation would be simplified if high-index DAE solvers become available that have built-in options for mixed continuous-discrete simulation.
Data Availability
All software and data will be made available upon publication of the manuscript.
@@ Acknowledgments
Appendix A Control Vector Parameterization
A.1 Piecewise constant and linear controls
Piecewise constant and linear controls are most common in optimal control (Nolasco et al., 2021). This section justifies the use of piecewise linear controls in Problems 2 and 3 for the optimization-based approaches, in comparison to piecewise constant controls. This comparison considers the most complex problem (Problem 3), with opt_Ipopt as the solver.
Table 4: Comparison between the errors $||e||_{2}$ resulting from piecewise constant and linear control vector parameterization for different values of control intervals $n_{c}$.

### v2

@@ 4.2.1 Problem description and formulation
Instead of using the maximum temperature as in Problem 1, some past studies considered cases in which the freezing and thawing rates are controlled, i.e., the rate of temperature change is kept constant (Seki and Mazur, 2008; Jang et al., 2017; Baboo et al., 2019; Bojic et al., 2021). As such, this problem focuses on controlling the rate of temperature change during thawing by manipulating the heater temperature. The definition of the average temperature is as given in (10).
For a fixed heater temperature, the rate of temperature change is not constant (Figs. 2DE). Thus, the heater temperature needs to be manipulated. Consider the optimal control problem
@@
where $d\Theta_{\textrm{avg}}/d\tau$ is the rate of temperature change and $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ is the target value (setpoint), set to 0.04 for demonstration in this problem. For the optimization-based approaches, the control vector is parameterized using a piecewise linear function, with the number of control intervals $n_{c}=16$ (Appendix B).
@@ 4.2.2 DAE Reformulation
For the simulation-based technique, first analyze the objective function in (17). This objective function is minimized when the rate of temperature change is equal to the setpoint. Replacing the objective function in (17) with the algebraic equation $d\Theta_{\textrm{avg}}/d\tau=(d\Theta_{\textrm{avg}}/d\tau)_{\textrm{sp}}$ results in the system of DAEs
@@
where $\Theta_{b}$ is now treated as a state. If the liquid domain used for calculating the average temperature includes the boundary point, the differential index of (18) is 1. If the boundary point is excluded, the index is 2. In any cases, (18) can be solved using GEKKO as explained in Section 3.3.
@@ 4.2.3 Solution comparison
where $(d\Theta_{\textrm{avg}}/d\tau)_{k}$ is the rate of temperature change resulting from solving (17) or (18) evaluated at each time point $k$ in the time span [$\tau_{0},\tau_{f}$] and $n_{k}$ is the number of sampling time points. A small value of $||e||_{2}$ corresponds to the rate of temperature change being close to the target value, indicating that the heater temperature is optimal; i.e., the solution method is accurate.
The optimal solutions to Problem 2 obtained from the optimization- and simulation-based approaches are illustrated in Fig. 3. Most solution techniques predict the same heater temperature profile except the methods with fmincon, which predict a slightly higher temperature (Fig. 3A). As time progresses, the heater temperature increases to compensate for a reduction in the temperature difference (driving force) between the heater and product. The rate of temperature change can be controlled nearly perfectly at 0.04 for IPOPT, CasADi, GEKKO, and the simulation-based method, while this value fluctuates up to about 0.046 for fmincon, implying that fmincon fails to converge to the optimal solution in this case (Fig. 3B). It is also evident that the average temperature increases linearly at the rate of 0.04 in all cases except for fmincon (Fig. 3C).
Figure 3: Trajectories of the optimal (A) heater temperature, (B) rate of temperature change, and (C) average temperature simulated via the optimization- and simulation-based approaches for Problem 2.
@@
The simulation-based approach is by far the fastest solution method, accelerating the computation by 40$\times$ to 90$\times$ compared to the typical solvers IPOPT and fmincon (Table 2). Without parallel computing, optimization with IPOPT is fastest, followed by GEKKO, fmincon, and CasADi. Parallel computing can significantly reduce the computation time for fmincon. With the direct multiple shooting method (opt_mCasADi), the number of states resulting from spatial discretization is large. For example, with $n=20$, there are 41 states for each control interval and 656 states for $n_{c}=16$. This large nonlinear optimization is computationally expensive to solve. In terms of accuracy, fmincon produces the largest error, which is about 30–40 fold higher than for the other techniques. This indicates that the solution obtained from fmincon is not optimal, agreeing with the plots shown in Fig. 3. Although both fmincon and IPOPT employ the interior point methods, the detailed implementations (e.g., gradient/hessian calculation, scaling) are not identical, which could result in different accuracy and performance (see detailed discussion and comparison of these solvers in Rojas-Labanda and Stolpe (2015)). The simulation-based and other optimization-based methods except fmincon have similar accuracy. In this problem, the solution is now known a priori as in Problem 1, but the simulation-based approach is still much faster than the optimization-based techniques while giving the same level of accuracy.
@@ 4.3.1 Problem description and formulation
One of the primary interests in a Stefan problem is the moving solid-liquid interface. The literature has studied and shown the importance of interface tracking and optimization in various applications (Hill and Wu, 1994; Brezina et al., 2018; Srisuma et al., 2023b, for example). This problem focuses on controlling the interface velocity, i.e., the evolution of a melting/freezing process.
@@
where $dS/d\tau$ is the interface velocity and $\left(dS/d\tau\right)_{\mathrm{sp}}$ is the target velocity, set to $-0.1$ for demonstration in this problem. In Problem 1 (Figs. 2BC), the interface velocity is not constant when the heater temperature is fixed. In this problem, the aim is to control the interface velocity to be constant at $(dS/d\tau)_{\mathrm{sp}}$ by manipulating the heater temperature throughout the process as formulated in (20). For the optimization-based approaches, the control vector is parameterized using a piecewise linear function, with the number of control intervals $n_{c}=12$ (see Appendix B).
@@ 4.3.2 DAE Reformulation
To apply the simulation-based technique, a similar approach used in Problem 2 can be used, but here we instead enforce the interface velocity by modifying the interface equation (6) to be equal to $(dS/d\tau)_{\textrm{sp}}$. Consequently, (20) is reformulated as the DAE system
@@
The differential index of (21) is $n$, which is the number of grid points resulting from spatial discretization. Therefore, if the spatial discretization of the domain is made finer to increase numerical accuracy, the differential index of (21) also increases. In this case, $n=20$, and hence (21) becomes an index-20 DAE system, making the problem even more complicated and difficult to solve than Problems 1 and 2.
To explore the possibility of using the index reduction technique, MATLAB’s reduceDAEIndex was used to reduce the differential index of (21). However, the index reduction process was not complete even after 20 hours of simulation. Obviously, this technique is not feasible for our problem, as explained before in Section 3.3. Therefore, GEKKO’s DAE solver is needed.
@@ 4.3.3 Solution comparison
The optimal solutions to Problem 3 are shown in Fig. 4. The optimal heater temperature predicted by each technique is nearly identical except that there is a large drop near the end of the process for fmincon (Fig. 4A). The interface velocity can be controlled at about $-0.1$ throughout the process for IPOPT, CasADi, GEKKO, and the simulation-based technique, while some fluctuation is observed for fmincon (Fig. 4B). This indicates that fmincon is less accurate than the others approaches, which is a similar trend observed before in Problem 2. With the interface velocity controlled, the interface position recedes linearly at the rate of 0.1 (Fig. 4C).
Figure 4: Trajectories of the optimal (A) heater temperature, (B) interface velocity, and (C) interface position via the optimization- and simulation-based approaches for Problem 3.
For the computational performance, the simulation-based technique is about 60$\times$ to 900$\times$ faster than all of the optimization-based approaches (Table 3), despite an increase in the differential index. IPOPT, GEKKO, and CasADi are comparable for the direct single shooting method, while the multiple shooting method is much slower. Parallel computing can significantly accelerate fmincon but is still much slower than the simulation-based approach. In terms of accuracy, the error of the simulation-based approach is about $2$$\times$$10^{-5}$, which is similar to those of the optimization-based approaches except for fmincon. With a relatively large error, fmincon is the least accurate method in this problem.
@@
The simulation-based approach, however, has one limitation associated with the DAE reformulation. As described in Section 3.3, the DAE reformulation is guaranteed to be optimal when at least one of the inequality constraints, bounds, or algebraic equations resulting from reformulating the objective function is active. This is typically the case when the objective function of the original optimal control problem can be minimized by solving each subproblem resulting from control vector parameterization individually – in other words, local minimization at every control/time interval (subproblem) is equivalent to global minimization of the whole process (original problem). For example, in Problem 1, if the thawing time of every control interval is minimized, the total thawing time is automatically minimized. In Problems 2 and 3, if the rate of temperature change and interface velocity are controlled locally for each control interval, the whole process is also controlled. These objective functions are common in many applications, for which the simulation-based technique can be applied, including lithium-ion batteries of various chemistries (Berliner et al., 2022; Galuppini et al., 2023; Matschek et al., 2023) and microwave lyophilization (Srisuma et al., 2023a). An example of a system in which the method is not optimal is a reactor with multiple chemical reactions, with the goal of maximizing the amount of desired product at the end of the process. Maximizing the amount of product during the first control interval could concurrently increase the amount of some byproducts that can degrade the desired product later. In this situation, local optimization is not equivalent to global optimization. As such, an algebraic equation/constraint may not be active, so an optimization-based technique is needed. Although the simulation-based approach does not give an optimal solution to such problems, it can be used to provide an initialization to the optimization solver.
@@ 4.4 Problem 4: Sensitivity analysis
In Problems 2 and 3, the simulation-based approach is tested with one set of values of $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ and $(dS/d\tau)_{\textrm{sp}}$, which are 0.04 and $-0.1$, respectively. To demonstrate the robustness of our framework, this section conducts a sensitivity analysis by varying $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ and $(dS/d\tau)_{\textrm{sp}}$ and employs the simulation-based approach to solve the problems.
The simulation-based approach provides accurate solutions irrespective of the values of $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ and $(dS/d\tau)_{\textrm{sp}}$ (Figs. 5 and 6). The rate of temperature change and interface velocity are at the setpoints except for $(dS/d\tau)_{\textrm{sp}}=-0.15$. An interface velocity of $-0.15$ is too fast for the given heater temperature, so the upper bound is active, and that the heater temperature cannot be increased further to achieve the target velocity. A target velocity that is larger (smaller in magnitude) than the peak of Fig. 2C, which about $-0.12$, will never violate the bound (see Appendix C for a formal proof).
For cases where the bound is active, a simulation-switching technique can be used to transition between active constraints, resulting in a hybrid discrete/continuous dynamic simulation (Feehery and Barton, 1998; Barton et al., 2000; Berliner et al., 2022). First, initialize a DAE simulation as usual. Next, terminate the current simulation when the bound on any control vector is active and then initialize a new DAE simulation with that active bound/constraint. Perform this switching whenever there is a change in the active constraints, which results in a mixed continuous-discrete DAE system instead of a pure continuous system. The complexity of this procedure depends on the choice of a DAE solver. For example, MATLAB’s and Julia’s DAE solvers have a built-in function to track all variables and terminate a simulation when the specified condition (aka event) is met, which facilitates the implementation of such switching. The switching technique can also be used to handle cases where the control trajectory is discontinuous (see examples in Berliner et al. (2022); Srisuma et al. (2023a)). GEKKO’s DAE solvers, on the other hand, do not have a built-in option for handling switches, and so this process needs to be executed manually. This switching technique can also be extended to handle path constraints.
Figure 5: Trajectories of the optimal (A) heater temperature, (B) rate of temperature change, and (C) average temperature simulated via the simulation-based approach at four different values of $(d\Theta_{\textrm{avg}}/d\tau)_{\mathrm{sp}}$ for Problem 4.
Figure 6: Trajectories of the optimal (A) heater temperature, (B) interface velocity, and (C) interface position simulated via the simulation-based approach at four different values of $(dS/d\tau)_{\textrm{sp}}$ for Problem 4.
@@ 5 Conclusion
Our proposed framework, the simulation-based approach, is demonstrated and benchmarked against a variety of optimization-based approaches – IPOPT, fmincon, CasADi, and GEKKO – for case studies involving the optimal control of a Stefan problem for cell thawing. The case studies include the minimization of the thawing time, the control of the average temperature, and the control of the interface velocity, covering both index-1 and high-index DAE systems. The simulation-based approach is faster than every optimization-based method by more than an order of magnitude while giving similar/better accuracy in all cases. The current simulation-based approach, however, is limited to problems where at least one of the inequality constraints, bounds, or algebraic equations resulting from reformulating the objective function is active.
Some future directions of interest are to (1) generalize the approach for more complicated problems, objective functions, and constraints that arise in multiple fields and (2) improve the computational performance with more efficient solvers and implementation.
@@ Acknowledgments
Appendix A Solver Options
There are a large number of options for optimization, ODE, and DAE solvers, which could lead to different accuracy and computational performance. This section describes and justifies the choice of important solver options used in our simulations. The default values are used for options that are not listed here.
For all case studies, the selected DAE solvers are MATLAB’s ode15s for index-1 DAEs and GEKKO’s DAE solver (Python) for high-index systems. The differential index of DAEs can be checked using MATLAB’s reduceDAEIndex. The wall-clock times (aka wall times, clock times) can be measured using the tic and toc functions, with each simulation repeated at least 10 times until the standard deviation of the measured wall times is smaller than 5% for consistent results. The initial guess of $\Theta_{b}$ is set to 0.5, the average dimensionless temperature, so that every solution method is initialized and compared on the same basis.
For the optimization-based approaches, the optimality tolerance is set to $10^{-7}$ for all optimization solvers, including IPOPT, fmincon, CasADi, and GEKKO. The integration tolerance for ode15s is $10^{-10}$. Using higher values for the optimality and integration tolerances could lead to inaccurate solutions, while using tighter values does not improve the accuracy significantly. For the objective functions (17) and (20) in Problems 2 and 3, the derivatives are approximated using a finite difference scheme, whereas the integrals can be approximated using a Riemann sum. The time step $\Delta\tau$ is set to 0.05.
While ode15s uses adaptive time-stepping, this option is not available in GEKKO; i.e., users manually specify a vector of time points for ODE integration. Hence, there is no ideal comparison between GEKKO’s ODE solvers and MATLAB’s ode15s. The fixed time step $\Delta\tau=0.05$ is specified for ODE integration in GEKKO, with five collocation points (INODES = 5) for each time interval. This time step is chosen to be consistent with the value used for a finite difference and Riemann sum approximation mentioned in the previous paragraph, while the number of collocation points is selected such that that the temperature and interface position simulated by GEKKO’s ODE solver have the same level of accuracy as those simulated by ode15s.
For the simulation-based approach, ode15s is used for an index-1 system (Problem 1), with the exact same solver options as mentioned above. GEKKO’s DAE solver is used for high-index systems (Problems 2 and 3). As a reformulated DAE system is different from the original ODE system, the time step needs to be adjusted accordingly. The time step is selected such that the accuracy of the simulation-based solution measured by $||e||_{2}$ is on the same order of magnitude as that of the optimization-based solution, giving $\Delta\tau=0.12$ and $\Delta\tau=0.37$ for Problems 2 and 3, respectively.
The accuracy of the optimization- and simulation-based approaches is measured using $||e||_{2}$ defined by (19) and (22). To calculate $||e||_{2}$, the derivatives are approximated using a finite difference scheme with $\Delta\tau=0.05$.
Appendix B Control Vector Parameterization
This appendix describes the choice of control vector parameterization used for the optimization-based approaches.
B.1 Piecewise constant and linear controls
Piecewise constant and linear controls are most common in optimal control (Nolasco et al., 2021). This section justifies the use of piecewise linear controls in Problems 2 and 3 for the optimization-based approaches, in comparison to piecewise constant controls. This comparison considers the most complex problem, Problem 3, and uses opt_Ipopt as the solver.
Table 4: Comparison between the errors $||e||_{2}$ of piecewise constant and linear control vector parameterization for different values of control intervals $n_{c}$.